OpenStudy (anonymous):

(-4 × 6) + 8 ______________ -2

OpenStudy (anonymous):

\[\frac{ (-4*6)+8 }{ -2 } = \frac{ -24+8 }{ -2 } = \frac{ -16 }{ -2 } = 8\] P.S: \[\frac{ - }{ - } = + \]